I really appreciate the help and have attempted to solve the problem below.
Did I do this correctly?
Please help...
-2(4x+1)=24
my answrer: -13/16
Any Suggestions?

Answer by mananth(16946)

-2(4x+1)=24
..
-8x-2 = 24
add +2
-8x-2+2 = 24+2
-8x = 26
divide by -8
x= 26/-8
x= -13/4
...
Plug value in the equation to check
-2(4*(-13/4)+1)
-2(-12)
=24
...
m.ananth@hotmail.ca